# One-Shot Real-World Demonstration Synthesis for Scalable Bimanual Manipulation

### v1

@@ 4.1 Deconstruction of One-Shot Teaching
From Blocks to Atomic Execution Primitives (AEPs). To enable modular adaptation, each block $\mathcal{B}_{i}$ is further refined into a minimal motion unit AEP, where an arm undergoes a salient state transition (e.g., gripper closure, moving or pose shifts). For arm $A\!\in\!\{L,R\}$, an AEP is defined as:
@@
$\centering\texttt{AEP}_{A}=\left\{\textbf{s}_{t}^{A}\rightarrow\textbf{s}_{t+\Delta t}^{A}\;|\;\left\|\textbf{s}_{t+\Delta t}^{A}-\textbf{s}_{t}^{A}\right\|_{2}\geq\gamma,\Delta t\leq T_{\textrm{max}},\texttt{CollisionFree}(\textbf{s}_{[t,\Delta t]}^{A})\right\},\@add@centering$ (6)
@@
where $\gamma$ is a distance threshold, $T_{\textrm{max}}$ limits execution duration, and $\texttt{CollisionFree}(\cdot)$ enforces no collisions between the arm and objects/scene (optimized via subsequent forward kinematics). Unlike keyframe-based segmentation james2022coarse; shridhar2023perceiver; ma2024hierarchical, AEPs ignore acceleration profiles (assuming quasi-static motions) and prioritize task-oriented transitions over temporal granularity.
@@
Semantic Categorization for Adaptive Synthesis. The final step categorizes refined blocks into invariant and variable types. The invariant blocks encode task-semantic primitives (e.g., screwing, pressing) or general motions (e.g., lifting, transferring), which remain structurally consistent across variations. The variable blocks, such as object-centric goal-conditioned grasping, adapt to instance-level geometric variations including object poses and shapes through:
@@
$\centering\mathcal{B}^{\prime}_{i}=\mathcal{B}_{i}\circ\Phi(g,\textbf{o}_{\textrm{novel}}),\@add@centering$ (7)
@@
where $\Phi(\cdot)$ is a vision-driven adapter, $g$ is the task goal, and $\textbf{o}_{\textrm{novel}}$ is the novel observation. This structured decomposition enables selective diversification, allowing variable blocks to adapt to new scenarios while preserving task fidelity in invariant blocks.
4.2 Vision-based Initial Frame Alignment
@@
The variable blocks identified in Sec. 4.1 (e.g., initial grasping) require adapting to novel object poses and geometries. To achieve this, we propose a vision-driven adapter $\Phi(\cdot)$ that generalizes object-centric interactions in the given exemplar to new scenes. Given a novel observation $\textbf{o}_{\textrm{novel}}$, our method executes three steps: object perception, state estimation, and pose alignment, ensuring the initial robotic grasp aligns with the task goal $g$. Illustrations are shown in Fig. 3.
@@
Figure 3: Illustrations of the initial frame alignment stage applied to tasks pouring (left and middle) and reorient (right). It shows that we can automatically adjust the grasp pose after the position, orientation and shape of the manipulated object changes.
@@
Object Perception. We first detect and segment the target object in $\textbf{o}_{\textrm{novel}}$ using an open-vocabulary detector (e.g., YOLO-World cheng2024yolo or YOLOE wang2025yoloe) or vision foundation models (e.g., Florence2 xiao2024florence with SAM2 ravi2024sam for robustness to rare categories). In practice, we prioritize foundation models to avoid detection failures. Now let $M\!\subseteq\!\textbf{o}_{\textrm{novel}}$ denote the obtained object’s binary mask.
@@
State Estimation. Instead of uitlizing category-level CAD-based 6D pose estimation models (e.g., FoundationPose wen2024foundationpose), we estimate the instance-level 6D object pose $\mathbf{P}_{\textrm{obj}}\!\in\!SE(3)$ using geometry-aware processing. Sepcifically, we adopt classic image moments chaumette2004image; kotoulas2007accurate to calculate the object mask centroid c and extract principal axes R:
@@
where $\textrm{d}(u,v)$ is the depth value. PCA is used to fit 3D points of $M$ to determine orientation $\textbf{R}\!\in\!SO(3)$. This yields $\mathbf{P}_{\textrm{obj}}\!=\!(\textbf{R},\textbf{c})$, robust to arbitrary object states (e.g., fallen, inverted).
Pose Alignment. Let $\mathbf{P}_{\textrm{demo}}$ denote the object pose in the given demonstration. We compute the rigid transformation $\mathbf{T}\!\in\!SE(3)$ that maps $\mathbf{P}_{\textrm{demo}}$ to $\mathbf{P}_{\textrm{obj}}$. This transformation is applied to the initial grasp pose $\mathbf{P}_{\textrm{grasp,demo}}$ in the variable block $\mathcal{B}_{i}$, yielding the adapted grasp pose:
$\centering\mathbf{P}_{\textrm{grasp,novel}}=\mathbf{T}\cdot\mathbf{P}_{\textrm{grasp,demo}}=(\mathbf{P}_{\textrm{obj}}\cdot\mathbf{P}^{-1}_{\textrm{demo}})\cdot\mathbf{P}_{\textrm{grasp,demo}}.\@add@centering$ (9)
The robot arm then executes a collision-free motion to reach $\mathbf{P}_{\textrm{grasp,novel}}$, ensuring goal-conditioned grasping without relying on task-irrelevant dense grasp proposals produced via 6-DoF grasp pose detectors fang2020graspnet; fang2023anygrasp. By integrating with the decomposition stage, the adapted grasp pose directly modifies the variable block $\mathcal{B}_{i}$ into $\mathcal{B}^{\prime}_{i}$, enabling synthesis of new trajectories.
4.3 Trajectory Modulation and Optimization
After deconstructing the demonstration into blocks (Sec. 4.1) and adapting variable blocks via vision-based alignment (Sec. 4.2), the synthesized trajectory requires two critical refinements to ensure executability and robustness as below:
Collision-Aware Validation. Each adapted block undergoes kinematic feasibility checks. First, we validate the reachability of target poses $\mathbf{P}_{\textrm{grasp,novel}}$ using Inverse Kinematics (IK), ensuring the robot joint limits are satisfied. If the result returned is singular or unreachable (rarely happens), we will delete this sample. Second, a motion planner chitta2012moveit; schulman2014motion verifies potential collision between arms and scene objects by solving for collision-free paths between the start and end states in each block. This simplified two-point constraint reduces computational overhead while preserving safety, assuming the trajectory of original given demonstration is collision-free.
Instance-Level Motion Adaptation. To handle geometric variations across object instances with differing shapes and sizes, we adjust motion primitives in variable blocks based on the object 3D bounding box dimensions. Let $\Delta l$, $\Delta w$ and $\Delta h$ denote the length, width and height differences between the novel object bounding box $\textbf{b}_{\textrm{novel}}$ and the original $\textbf{b}_{\textrm{demo}}$. Motion endpoints (e.g., grasp or release positions) are offset by:
@@ 4.3 Trajectory Modulation and Optimization
where $\lambda$ scales the adjustment (empirically set to $0.8\sim 1.0$), and $A\!\in\!\{L,R\}$. For irregular shapes, depth-aware masks $M$ extracted in Sec. 4.2 can approximate volumetric differences, or minimal human input provides precise measurements.
@@
This hierarchical optimization stage operates at the block level rather than full trajectories, and enables efficient scaling while preserving the task semantics encoded in invariant blocks. Finally, validated and adapted blocks are recombined into synthetic trajectories $\tau^{\prime}\!=\!\{\tau_{\textrm{inv}},\{\mathcal{B}^{\prime}_{i}\}\}$, populating diverse and physically feasible demonstrations $\mathcal{D}_{syn}$.
5 Experiments
@@
Our experiments address three core questions. $\mathsf{Q}1$: Is BiDemoSyn truly efficient and user-friendly? $\mathsf{Q}2$: Does BiDemoSyn enable scalable visuomotor imitation learning? $\mathsf{Q}3$: Does synthetic demonstrations generalize to spatial and object variations?
5.1 Experimental Setup and Protocol
Tasks. We evaluate on six bimanual manipulation tasks (refer Fig. 9) requiring contact-rich coordination: plugpen, inserting, unscrew, pouring, pressing and reorient. These tasks cover diverse primitive skills (e.g., grasping, rotating and handover), and involve both rigid and articulated objects. The one-shot demonstration for each task is collected via kinesthetic teaching. For the platform (refer Fig. 8), two fixed-base arms equipped with parallel grippers are used. Scene perception is provided by a stereo camera capturing binocular images. More details of all tasks and the corresponding data collection process are in Appendix Sec. A.
Policies & Baselines. For a comprehensive comparison, baselines contain two categories. One category is purely for data collection including point cloud editing (DemoGen xue2025demogen), real robot auto-rollout (YOTO zhou2025you), and human drag teaching (close to Teleoperation). Generally, the quality of collected demonstrations by these baselines is better in turn, but the cost is more time-consuming. After preparing the training data, we adapt three advanced visuomotor policies (DP chi2023diffusion, DP3 ze2024dp3 and EquiBot yang2024equibot) to bimanual settings by modifying their modeling spaces to dual-arm actions (refer Sec. 3). Observation inputs are RGB-only images or segmented 3D point clouds of task-relevant objects. Final trained policies operate in the open-loop discrete keyposes prediction to align with our synthesized demonstration format. The another category is directly for bimanual manipulation without retraining including the zero-shot ReKep huang2024rekep, an advanced ReKep+ with oracle-level grasp labels at the beginning, and one-shot ODIL wang2025one.
@@ 5.1 Experimental Setup and Protocol
Figure 4: (A) The data collection efficiency comparison of different baselines and our BiDemoSyn, and (B) the generated data quality illustration of DemoGen. Although DemoGen has the highest synthesis efficiency, it cannot avoid visual artifacts caused by perspective transformation, so its data quality is the lowest. Our method can achieve a balance between the speed and quality.
5.2 Comparison and Presentation of Results
($\mathsf{A}1$): The efficiency and usability of BiDemoSyn have obvious advantages over baselines. Without bells and whistles, our BiDemoSyn demonstrates superior efficiency and usability as illustrated in Fig. 4. It requires about 5 seconds per demonstration to synthesize new trajectories, where primarily involving manual object repositioning and visual verification of initial frame alignment, followed by the trajectory optimization. In contrast, DemoGen generates trajectories in less then 1 second via scripted edits but suffers from perspective distortion artifacts, degrading data quality. The specific qualitative effect can be compared with Fig. 4A and our results in Fig. 2 and Fig. 9. By relying on manual alignment and auto-replay, YOTO takes 42 seconds per demo which is equivalent to real robot execution time. While, teleoperation demands 91 seconds per demo for near-perfect but labor-intensive collection. These results highlight BiDemoSyn’s unique balance of speed (automated optimization) and reliability (artifact-free synthesis), making it the only method scalable to thousands of demonstrations without compromising real-world fidelity.
@@ 5.2 Comparison and Presentation of Results
Table 1: Quantitative comparison results of success rates for different methods under in-distribution and out-of-distribution evaluations. The # means the BiDemoSyn training data size.
  in-distribution evaluations out-of-distribution evaluations
Policy
(Input
Mode)
Method
@@
Average
Success
Rate
@@
Average
Success
Rate
@@
Training
-Free
ReKep 15/30 13/30 12/30 15/30 12/30 09/30 42.2% 12/30 10/30 10/30 09/30 09/30 06/30 31.1%
@@
Figure 5: Comparison between training data scale and the success rate. Less sized data is randomly sampled out of the total dataset at the task level. DP3 is chosen as the visuomotor policy.
($\mathsf{A}2$): Demonstrations obtained via BiDemoSyn can support scalable imitation learning. We applied BiDemoSyn for efficient collection of thousands demonstrations per task (Fig. 9 right), densely covering the workspace with instance-level object diversity, thereby providing abundant training data for imitation learning. For two reproduced baselines, DemoGen xue2025demogen matches our data scale but suffers from visual artifacts, while YOTO zhou2025you collects only 1/10 the data per task for being limited by its time-consuming replay mechanism. After obtaining adequate data, we trained three advanced visuomotor policies DP chi2023diffusion, DP3 ze2024dp3 and EquiBot yang2024equibot. As shown in Tab. 1 left, BiDemoSyn achieves superior in-distribution (ID) success rates across all six tasks, outperforming both baselines regardless of the policy architecture. This confirms that BiDemoSyn’s artifact-free synthesis and efficient data scaling reliably support the scaling laws of imitation learning. Besides, it is unsurprising that policies trained with demonstrations consistently outperform the two training-free baselines ReKephuang2024rekep and ODIL wang2025one. To further quantify scalability, we evaluated training data size-performance relationships on all six tasks of all methods. As shown in Fig. 5, BiDemoSyn-trained policies exhibit stronger scaling trends than DemoGen (e.g., 60% vs. 39% at 1106 demos), aligning with community observations that data quality and quantity jointly drive visuomotor policy performance liu2025rdt; black2024pi0; pertsch2025fast; lin2025data.
@@
Figure 6: Analysis of spatial generalization, including variations in (A) position sampling density and (B) orientation sampling diversity. DP3 is chosen as the visuomotor policy.
($\mathsf{A}3$): Policies trained on BiDemoSyn data can achieve better generalization to unseen variations. To evaluate the generalization to novel instances, we conduct controlled tests across all six tasks. For each task, we exclude demonstrations associated with one or a pair of randomly selected objects (e.g., an arbitrary bottle in unscrew or a bottle-cup pair in pouring) from the training set, resulting in training sizes of 2916/3888/1008/972/324/756 demos for six tasks, respectively. These excluded objects are reintroduced exclusively during real-world testing, with identical data splits applied to DemoGen and YOTO baselines for fair comparison. As shown in Tab. 1 right, policies trained on BiDemoSyn data consistently achieve higher success rates compared to all baselines in out-of-distribution (OOD) settings. While all methods exhibit performance drops relative to ID evaluations, BiDemoSyn’s superior OOD robustness highlights its ability to capture task-invariant features through vision-aligned synthesis. To further analyze spatial generalization, we vary the density of positional and orientation coverage in training data: for unscrew, we synthesize trajectories with sparse-to-dense workspace coverage (Fig. 6A), and for reorient, we incrementally increase the diversity of object orientations (Fig. 6B). To exclude the effect of data size differences, the total number of train-set is always augmented to a standard size (e.g., #1152 and #1008) via the DemoGen to edit some observations at small distances. Finally, policies trained on BiDemoSyn data with dense spatial coverage (e.g., $10\times 10$ workspace coverage in unscrew) achieve 76.7% success on unseen positions, compared to 53.3% for sparse coverage ($6\times 6$ grid cells). Similarly, orientation diversity in reorient improves generalization from 37.3% (lower diversity with #Ornt.=3) to 73.3% (higher diversity with #Ornt.=7). These results confirm that BiDemoSyn’s ability to cheaply synthesize positionally and orientationally diverse demos directly translates to enhanced policy generalization, aligning with empirical insights that broad data coverage mitigates spatial distribution shifts.

### v2

@@ IV-A Deconstruction of One-Shot Teaching
IV-A1 From Blocks to Atomic Execution Primitives (AEPs)
@@
To enable modular adaptation, each block $\mathcal{B}_{i}$ is further refined into a minimal motion unit AEP, where an arm undergoes a salient state transition (e.g., gripper closure, moving or pose shifts). For arm $A\!\in\!\{L,R\}$, an AEP is defined as:
@@
$\centering\texttt{AEP}_{A}=$ $\{\textbf{s}_{t}^{A}\rightarrow\textbf{s}_{t+\Delta t}^{A}\;|\;\left\|\textbf{s}_{t+\Delta t}^{A}-\textbf{s}_{t}^{A}\right\|_{2}\geq\gamma,$ (6)
$\Delta t\leq T_{\textrm{\scriptsize max}},\texttt{CollisionFree}(\textbf{s}_{[t,\Delta t]}^{A})\},$
@@
where $\gamma$ is a distance threshold, $T_{\textrm{\scriptsize max}}$ limits execution duration, and $\texttt{CollisionFree}(\cdot)$ enforces no collisions between the arm and objects/scene (optimized via forward kinematics). Unlike keyframe-based segments [39, 78, 61, 107], AEPs ignore acceleration profiles (assuming quasi-static motions) and prioritize task-oriented transitions over temporal granularity.
IV-A2 Semantic Categorization for Adaptive Synthesis
@@
The final step categorizes refined blocks into invariant and variable types. The invariant blocks encode task-semantic primitives (e.g., screwing, pressing) or general motions (e.g., lifting, transferring), which remain structurally consistent across variations. The variable blocks, such as object-centric goal-conditioned grasping, adapt to instance-level geometric variations including object poses and shapes through:
@@
$\centering\mathcal{B}^{\prime}_{i}=\mathcal{B}_{i}\circ\Phi(g,\textbf{o}_{\textrm{\scriptsize novel}}),\@add@centering$ (7)
@@
where $\Phi(\cdot)$ is a visual adapter, $g$ is the task goal, and $\textbf{o}_{\textrm{\scriptsize novel}}$ is the novel observation. This structured decomposition enables selective diversification, allowing variable blocks to adapt to new scenes while preserving task fidelity in invariant blocks.
IV-B Vision-based Initial Frame Alignment
@@
The variable blocks identified in Sec. IV-A require adapting to novel object poses and geometries. To achieve this, we propose a vision-driven adapter $\Phi(\cdot)$ that generalizes object-centric interactions in the given exemplar to new scenes. Given a novel observation $\textbf{o}_{\textrm{\scriptsize novel}}$, our method executes three steps: object perception, state estimation, and pose alignment, ensuring the initial robotic grasp aligns with the task goal $g$. Illustrations are shown in Fig. 2.
@@
Figure 2: Illustrations of the initial frame alignment applied to tasks pouring (left and middle) and reorient (right). It shows that we can automatically adjust the grasp pose after the position, orientation and shape of the manipulated object changes.
IV-B1 Object Perception
@@
We first detect and segment the target object in $\textbf{o}_{\textrm{\scriptsize novel}}$ using an open-vocabulary detector (e.g., YOLO-World [14] or YOLOE [81]) or vision foundation models (e.g., Florence2 [89] with SAM2 [74] for robustness to rare categories). In practice, we prioritize foundation models to avoid detection failures. Now let $M\!\subseteq\!\textbf{o}_{\textrm{\scriptsize novel}}$ denote the obtained object’s binary mask.
IV-B2 State Estimation
@@
Instead of uitlizing category-level CAD-based 6D pose estimation models (e.g., FoundationPose [86]), we estimate the instance-level 6D object pose $\mathbf{P}_{\textrm{\scriptsize obj}}\!\in\!SE(3)$ using geometry-aware processing. Sepcifically, we adopt classic image moments [8, 47] to calculate the object mask centroid c and extract principal axes R:
@@
where $\textrm{d}(u,v)$ is the depth value. PCA is used to fit 3D points of $M$ to determine orientation $\textbf{R}\!\in\!SO(3)$. This yields $\mathbf{P}_{\textrm{\scriptsize obj}}\!=\!(\textbf{R},\textbf{c})$, robust to arbitrary object states (e.g., fallen, inverted).
IV-B3 Pose Alignment
Let $\mathbf{P}_{\textrm{\scriptsize demo}}$ denote the object pose in the given demonstration. We compute the rigid transformation $\mathbf{T}\!\in\!SE(3)$ that maps $\mathbf{P}_{\textrm{\scriptsize demo}}$ to $\mathbf{P}_{\textrm{\scriptsize obj}}$. This transformation is applied to the initial grasp pose $\mathbf{P}_{\textrm{\scriptsize grasp,demo}}$ in the variable block $\mathcal{B}_{i}$, yielding the adapted grasp pose:
$\centering\mathbf{P}_{\textrm{\scriptsize grasp,novel}}=\mathbf{T}\cdot\mathbf{P}_{\textrm{\scriptsize grasp,demo}}=(\mathbf{P}_{\textrm{\scriptsize obj}}\cdot\mathbf{P}^{-1}_{\textrm{\scriptsize demo}})\cdot\mathbf{P}_{\textrm{\scriptsize grasp,demo}}.\@add@centering$ (9)
The robot arm then executes a collision-free motion trajectory to reach $\mathbf{P}_{\textrm{\scriptsize grasp,novel}}$, ensuring goal-conditioned grasping without relying on task-irrelevant dense grasp proposals produced via 6-DoF grasp pose detectors [24, 23]. By integrating with the proposed decomposition stage, the adapted grasp pose directly modifies the variable block $\mathcal{B}_{i}$ into $\mathcal{B}^{\prime}_{i}$, enabling synthesis of new trajectories.
IV-C Trajectory Modulation and Optimization
After deconstructing the provided demonstration into blocks (Sec. IV-A) and adapting variable blocks via vision-based alignment (Sec. IV-B), the synthesized trajectory still requires two critical refinements to ensure executability and robustness as expalined below:
IV-C1 Collision-Aware Validation
Each adapted block undergoes kinematic feasibility checks. First, we validate the reachability of target poses $\mathbf{P}_{\textrm{\scriptsize grasp,novel}}$ using Inverse Kinematics (IK), ensuring the robot joint limits are satisfied. If the result returned is singular or unreachable (rarely happens), we will delete this sample. Second, a motion planner [19, 75] verifies potential collision between arms and scene objects by solving for collision-free paths between the start and end states in each block. This simplified two-point constraint reduces computational overhead while preserving safety, assuming the trajectory of original given demonstration is collision-free.
IV-C2 Instance-Level Motion Adaptation
To handle geometric variations across object instances with differing shapes, we adjust motion primitives in variable blocks based on the object 3D bounding box dimensions. Let $\Delta l$, $\Delta w$ and $\Delta h$ denote the length, width and height differences between the novel object bounding box $\textbf{b}_{\textrm{\scriptsize novel}}$ and the original $\textbf{b}_{\textrm{\scriptsize demo}}$. Motion endpoints (e.g., grasp or release positions) are offset by:
@@ IV-C2 Instance-Level Motion Adaptation
where $\lambda$ scales the adjustment (empirically set to $0.8\sim 1.0$), and $A\!\in\!\{L,R\}$. For irregular shapes, depth-aware masks $M$ extracted in Sec. IV-B can approximate volumetric differences, or minimal human input provides precise measurements.
@@
This hierarchical optimization stage operates at the block level rather than full trajectories, and enables efficient scaling while preserving the task semantics encoded in invariant blocks. Finally, validated and adapted blocks are recombined into synthetic trajectories $\tau^{\prime}\!=\!\{\tau_{\textrm{\scriptsize inv}},\{\mathcal{B}^{\prime}_{i}\}\}$, populating diverse and physically feasible demonstrations $\mathcal{D}_{syn}$.
V Experiments
@@
Our experiments address five core questions. $\mathsf{Q}1$: Is BiDemoSyn truly efficient and user-friendly? $\mathsf{Q}2$: Does BiDemoSyn enable scalable visuomotor imitation learning? $\mathsf{Q}3$: Does synthetic demonstrations generalize to spatial and object variations? $\mathsf{Q}4$: Can BiDemoSyn seamlessly integrate with few-shot demonstrations to further improve robustness and data efficiency? $\mathsf{Q}5$: Do policies trained on BiDemoSyn data transfer zero-shot across different robot embodiments while maintaining high success rates?
@@
Figure 3: (A) The data collection efficiency comparison of different baselines and our BiDemoSyn, and (B) the generated data quality illustration of DemoGen. Although DemoGen has the highest synthesis efficiency, it cannot avoid visual artifacts caused by perspective transformation, so its data quality is the lowest. Our method can achieve a good speed-quality balance.
@@
TABLE I: Quantitative results of different methods under in-distribution (ID) and out-of-distribution (OOD) evaluations. The OOD means removing one or a pair of objects from the training set for testing purposes only. The # means the training size.
  in-distribution (ID) evaluations out-of-distribution (OOD) evaluations
Policy (Input Mode) Method
@@
Average Success Rate
@@
Average Success Rate
@@
  Training -Free (RGB) ReKep 15/30 13/30 12/30 15/30 12/30 09/30 42.2% 12/30 10/30 10/30 09/30 09/30 06/30 31.1%
@@
V-A Experimental Setup and Protocol
Tasks. We evaluate on six bimanual manipulation tasks requiring contact-rich coordination: plugpen, inserting, unscrew, pouring, pressing and reorient. These tasks cover diverse primitive skills (e.g., grasping, rotating and handover), and involve both rigid and articulated objects. The one-shot demonstration for each task is collected via kinesthetic teaching. For the platform, two fixed-base arms equipped with parallel grippers are used as the main workspace. An auxiliary humanoid dual-arm platform was also used to verify the cross-platform deployment capability of the learned policy. Scene perception is provided by a stereo camera capturing binocular images. More details of tasks, harwares and data collection are in Supplementary Materials.
Policies and Baselines. Compared baselines contain two categories. One category is purely for data collection including point cloud editing (DemoGen [93]), real robot auto-rollout (YOTO [107]), and human drag teaching (close to Teleoperation). Generally, the quality of collected demonstrations by these baselines is better in turn, but the cost is more time-consuming. After preparing the training data, we adapt three advanced visuomotor policies (DP [16], DP3 [98] and EquiBot [94]) to bimanual settings by modifying their modeling spaces to dual-arm actions (refer Sec. III). Observation inputs are RGB-only images or segmented 3D point clouds of task-relevant objects. Final trained policies operate in the open-loop discrete keyposes prediction to align with our synthesized demonstration format. The another category is directly for bimanual manipulation without retraining including the zero-shot ReKep [38], an advanced ReKep+ with oracle-level grasp labels at the beginning, and two one-shot imitation learning methods ODIL [84] and MAGIC [60].
V-B Comparison and Presentation of Results
@@
Figure 4: Comparison between training scale and success rate. Less sized data is randomly sampled out of the total dataset at task-level. DP3 is chosen as the visuomotor policy.
@@
Figure 5: Analysis of spatial generalization for variations in (A) position sampling density and (B) orientation sampling diversity. DP3 is chosen as the policy.